OpenStudy (anonymous):

Given the function f(x) = x^3 + 2x^2 − 3x − 5, what is the resulting function when f(x) is shifted to the right 1 unit? f(x + 1) = x^3 + 5x^2 + 4x − 5 f(x) + 1 = x^3 + 2x^2 − 3x − 4 f(x) − 1 = x^3 + 2x^2 − 3x − 6 f(x − 1) = x^3 − x^2 − 4x − 1

OpenStudy (anonymous):

@CaptainFluffy

OpenStudy (anonymous):

@paki

OpenStudy (captainfluffy):

umm ewww math

OpenStudy (captainfluffy):

well i can try

OpenStudy (captainfluffy):

ima take a slight guess...

OpenStudy (captainfluffy):

(x) = x3 + 2x2 - 3x - 5 Shifted to the right 1 unit f(x - 1) = (x - 1)^3 + 2(x - 1)^2 - 3(x - 1) - 5 = (x^3 - 3x^2 + 3x - 1) + 2(x^2 - 2x + 1) - 3(x - 1) - 5 = x^3 - 3x^2 + 2x^2 + 3x - 4x - 3x - 1 + 2 + 3 - 5 = x^3 - x^2 - 4x - 1 2) f(x) = x4 + 3x3 - 5x2 + 2x - 2 f(-1) = (-1)^4 + 3(-1)^3 - 5(-1)^2 + 2(-1) - 2 = -11 f(1) = (1)^4 + 3(1)^3 - 5(1)^2 + 2(1) - 2 = -1 average rate of change = [-11 - (-1)]/[1 - (-1)] = -5 3) f(x) = -3x^38 + 7x^3 - 12x + 13 f(-∞) = -3(-∞)^38 + 7(-∞)^3 - 12(-∞) + 13 = -∞ f(∞) = -3(∞)^38 + 7(∞)^3 - 12(∞) + 13 = -∞ Down on the left, down on the right 4) f(x) = 8x^3 - 3x^2 - 5x + 8 The degree of the first term 5) f(x) = x^4 - 7x^3 + 12x^2 + 4x - 12 From x = 1 to x = 2 .

OpenStudy (captainfluffy):

im mostly wrong tho

OpenStudy (captainfluffy):

so idk im not good at math at all

OpenStudy (anonymous):

Noo youre great thank you